# Applying Text Embedding Models for Efficient Analysis in Labeled Property Graphs

### v3

@@ Abstract
footnotetext: Preprint. Submitted to European Symposium on Artificial Neural Networks, Computational Intelligence and Machine Learning (ESANN) 2026.
@@ 1 Introduction
Labeled property graphs (LPGs) represent a flexible data model in which nodes and edges are not only connected through relationships but also enriched with key-value properties, most of which are textual [1]. This structure is well-suited to domains such as knowledge graphs, social networks, and institutional data, where descriptive attributes carry important semantic information. Despite this, most analytical approaches to LPGs focus primarily on structural information - such as connectivity patterns or edge types - while treating textual fields as flat labels or ignoring them altogether.
@@
In this work, we investigate how text embedding models can be used to enhance property graph analysis. Specifically, we apply them to two fundamental tasks: node classification and link prediction. In both cases, textual attributes are transformed into semantic vector representations using a state-of-the-art embedding model. These embeddings are then used as inputs to lightweight classifiers, enabling the graph to be analyzed through its language rather than just its structure.
@@ 2 Related work
Graph-based machine learning methods such as GCNs [2], GraphSAGE [3], and R-GCN [4] have shown strong performance in tasks like node classification and link prediction, but they largely ignore textual node and edge attributes. Meanwhile, pretrained text embedding models [5, 6] capture rich semantic information from text and are widely used in NLP tasks.
Recent work at the intersection of graphs and language includes combining structural and textual features in graphs [7, 8], but these methods often assume fixed schemas or require joint training. Property graphs, with their flexible attributes, remain underexplored in this context.
Our approach differs by integrating pretrained text embedding models into labeled property graph analysis without modifying the graph structure. This enables semantically informed classification and relation prediction using existing text fields as-is.
@@ 4 Solution
Our approach integrates pretrained text embedding models into the analysis of labeled property graphs, enabling semantic-aware workflows for two tasks: node classification and relation prediction. We employ the Qwen3-Embedding-0.6B model [16], a compact but high-performing text embedding model from Alibaba, which achieves state-of-the-art results on the Massive Text Embedding Benchmark (MTEB) [14, 15] despite its small size. This model allows us to efficiently embed textual node and edge properties into dense vector representations without modifying the structure of the graph or fine-tuning the model.
@@
We formulate two prediction tasks based on language-model embeddings of graph nodes. In both cases, the available textual and relational information for node is serialized into a coherent string and processed by the embedding model to produce a 1024-dimensional vector representation. Instances are randomly split into training and test sets, with 90% used for training and 10% reserved for evaluation. Standard classifiers - Random Forest, Logistic Regression, SGDClassifier, and Support Vector Machine (SVM) - are trained on the embeddings and evaluated on the held-out test set.
In the node label prediction task, the input string contains the textual properties of the node, combined into a single coherent description. The classifiers learn to map these embeddings to the corresponding node labels and can subsequently be applied to unlabeled nodes.
In the relation prediction task, the goal is to infer missing factual information (Figure 1). For each labeled instance, a specific relation of the source node is withheld (for example, the relation between a player and a team). The remaining relations of the source node, together with its connected neighbor nodes, are included in the textual description. The trained classifiers predict the correct target node, thereby recovering the withheld relation.
@@
Figure 1: Relation Prediction Setting: In this setting, a specific relation of the source node (e.g., between a Player and a Team in the WWC 2019 dataset) is withheld prior to embedding generation. The embedding model encodes the source node based on its remaining relations and neighboring nodes. The resulting embedding is then used in a downstream classification task to predict the correct target node, thereby recovering the withheld relation.
@@ 5 Results
Support Vector Machine 1.000 1.000 1.000 1.000
(a) WWC 2019
@@
Support Vector Machine 1.000 1.000 1.000 1.000
(b) Twitter Trolls
Classifier Accuracy Precision Recall F1 Score
Random Forest 0.822 0.823 0.822 0.821
Logistic Regression 0.926 0.926 0.926 0.926
SGDClassifier 0.928 0.930 0.928 0.929
Support Vector Machine 0.939 0.939 0.939 0.939
(c) Stack Overflow
@@ 6 Next Steps
Future work could focus on enhancing the integration of textual and structural information, for example by combining language-model embeddings with topology-aware graph encoders. Another promising direction is task-specific adaptation of the embedding space to capture domain-specific semantics and relation patterns. Scaling to large and dynamic graphs may require efficient incremental embedding updates and distributed processing pipelines. Finally, systematic evaluation across heterogeneous datasets would help to assess generalization and identify domain-dependent optimization strategies.
@@ 7 Conclusions
We presented an approach for enhancing labeled property graph analysis using pretrained text embedding models. By encoding textual properties into semantic vectors, we enable more informed classification and relation prediction without modifying the underlying graph structure.
This method is model-agnostic, lightweight, and compatible with existing property graph platforms. Our evaluation demonstrates its broad applicability across domains.
The results highlight the potential of treating textual node and edge attributes as semantically rich signals-supporting more expressive and effective analysis workflows within graph-based systems.
Acknowledgment
This manuscript acknowledges the use of ChatGPT [17], powered by the GPT-4o language model developed by OpenAI, to improve language clarity, refine sentence structure, and enhance overall writing precision.

### v4

# https://nask.pl/

@@ 1 Introduction
Graph analysis has become a central tool for modeling and reasoning over complex, interconnected data. In many practical settings, such graphs are enriched with substantial textual information attached to nodes and edges, capturing descriptions, metadata, and contextual details. Labeled property graphs (LPGs) provide a widely used representation for such data, particularly in domains such as knowledge graphs, social networks, and institutional data. Nevertheless, most analytical approaches primarily focus on structural aspects - such as connectivity patterns or edge types - while the semantic content of textual properties remains largely underutilized.
@@
We present a lightweight and general framework for integrating pretrained text embeddings into labeled property graph analysis, supporting semantic reasoning without modifying graph topology or retraining models. The approach is evaluated on two core tasks - node classification and link prediction - using semantic vector representations of textual properties as features for simple classifiers. This demonstrates that language-informed embeddings can enhance traditional graph analysis pipelines.
@@ 2 Related work
Graph representation learning.
A large body of work studies how to learn predictive models over graph-structured data. Early approaches relied on shallow structural embeddings such as DeepWalk and node2vec, which map nodes to vectors by optimizing objectives derived from random walks [19, 20]. More recent graph neural networks (GNNs) generalize message passing across neighborhoods and have become standard approach. Representative architectures include Graph Convolutional Networks (GCN) [2], GraphSAGE for inductive learning [3], Graph Attention Networks (GAT) [18], and relational variants such as R-GCN designed for multi-relational graphs [4]. While these methods typically excel at exploiting topology and typed edges, many practical graphs - including labeled property graphs - contain extensive attribute information (often textual) that is not always leveraged effectively by purely structural encoders.
Attributed and text-rich graphs.
A parallel line of work addresses graphs with node/edge attributes, including text. Common strategies concatenate handcrafted text features (e.g., TF-IDF) with structural features, or encode text with neural models and fuse it with GNN representations via concatenation, gating, or attention [21, 22]. There is also work on heterogeneous graphs (with multiple node/edge types) using specialized aggregation and type-aware transformations [23, 24]. However, these approaches often assume a relatively fixed schema, require end-to-end retraining, or rely on carefully curated feature engineering. In contrast, labeled property graphs are frequently schema-flexible: nodes and relationships can carry varying sets of properties, with missing fields and a mixture of short categorical fields and long free-form text, making uniform feature design and model training more challenging.
Language models for knowledge graphs and graph–text integration.
Several methods integrate pretrained language models with structured triples for knowledge-graph completion or link prediction. KG-BERT linearizes a triple (or a small neighborhood) and applies a Transformer to score plausibility [7]. KEPLER and related approaches combine language modeling with knowledge embedding objectives, typically requiring joint training and access to curated textual descriptions aligned with knowledge graph entities [8]. Other lines of work explore text-guided graph encoders, including Graph-BERT-style pretraining on subgraphs and textual signals, or hybrid architectures that encode both local topology and text [25]. Although effective in certain settings, these approaches are usually developed for RDF-like knowledge graphs with fixed predicates, or they impose a particular serialization/training pipeline. For native property graphs in operational graph databases, a lightweight method that can exploit available text without re-engineering the graph schema or training a large joint model remains practically valuable.
Positioning of our approach.
Compared to end-to-end graph–text fusion methods and KG-specific LM models, we adopt a deliberately lightweight strategy: we serialize available textual properties and local context, encode them with a pretrained text embedding model, and apply standard classifiers for node classification and relation prediction. This provides a simple interface between language semantics and LPG analytics: it requires no modification of the graph structure, no schema coupling, and no expensive retraining of graph encoders, while still enabling semantic-aware predictions when textual properties carry substantial signal.
3 Preliminaries
A labeled property graph (LPG) is a graph data model in which both vertices and edges are first-class entities enriched with semantic annotations [1]. Each node and relationship may carry one or more labels that define its conceptual type, as well as an arbitrary set of key–value properties used to store structured or unstructured attributes. Formally, an LPG can be defined as a directed multigraph $G=(V,E,L,P)$, where $V$ is the set of nodes, $E\subseteq V\times V$ is the set of edges, $L$ assigns one or more labels to nodes and edges, and $P$ maps each graph element to a set of properties. This flexible schema allows different nodes and relationships to exhibit heterogeneous structures and attribute sets. Labeled property graphs are widely adopted in practice due to their expressiveness and intuitive alignment with real-world entities and relations, making them well suited for domains such as knowledge graphs, social networks, and institutional data. However, this same flexibility introduces significant analytical challenges: the lack of a fixed schema, the heterogeneity of node and edge types, and the prevalence of high-dimensional and unstructured textual properties make it difficult for traditional graph analytics and learning methods to effectively capture semantic information.
@@ 5 Solution
Our approach integrates pretrained text embedding models into the analysis of labeled property graphs, enabling semantic-aware workflows for node- and relation-level prediction tasks. The framework is deliberately designed to be model-agnostic: any pretrained text embedding model that maps free-form text to fixed-length vector representations can be substituted without changes to the surrounding pipeline. In this study, we employ Qwen3-Embedding-0.6B [16] as a representative example, which achieves state-of-the-art performance on the Massive Text Embedding Benchmark [14, 15]. This choice enables efficient encoding of textual node and edge properties into dense vector representations without fine-tuning or modifying the underlying graph structure.
At a high level, the proposed pipeline treats textual properties as the primary carrier of semantic information within the graph. For each prediction instance, node-centric information is serialized into a structured textual representation, which is then encoded by the embedding model into a 1024-dimensional vector. These embeddings serve as general-purpose semantic features and are used as input to standard downstream classifiers. This design separates semantic representation learning from task-specific modeling, allowing the same embeddings to be reused across multiple tasks and classifiers.
@@
We consider two prediction tasks, both formulated over node embeddings. In both cases, instances are randomly split into training and test sets, with 90% of instances used for training and 10% reserved for evaluation. Downstream prediction is performed using a set of standard classifiers - Random Forest, Logistic Regression, SGDClassifier, and Support Vector Machine (SVM) - trained on the embedding vectors. This choice emphasizes the expressive power of the embeddings themselves rather than task-specific neural architectures, and facilitates transparent comparison across models.
In the node label prediction task, the objective is to predict the semantic label of a node based solely on its textual properties. The input text is constructed by concatenating all available textual properties of the node (excluding the label) into a single description, represented as a sequence of key–value pairs. This formulation reflects a common scenario in property graph analytics, where node types may be unknown or incomplete and must be inferred from descriptive attributes. The classifier learns a mapping from the embedding space to the label space and can subsequently be applied to unseen nodes without requiring access to graph structure or neighborhood information.
In the relation prediction task, the goal is to recover missing relational information by leveraging the textual context of a node and its local neighborhood (Figure 1). For each source node, one relation is withheld prior to embedding generation. The input text is then constructed from the source node’s textual properties, combined with the labels and properties of its remaining relations and directly connected neighbor nodes. This serialized representation captures both intrinsic node descriptions and immediate relational context in textual form. The classifier is trained to predict the correct target node associated with the withheld relation, effectively performing relation recovery based on semantic cues rather than explicit graph traversal.
@@
Figure 1: Relation Prediction Task Setting: In this setting, a specific relation of the source node (e.g., between a Player and a Team in the WWC 2019 dataset) is withheld prior to embedding generation. The embedding model encodes the source node based on its remaining relations and neighboring nodes. The resulting embedding is then used in a downstream classification task to predict the correct target node, effectively recovering the withheld relation.
@@ 6 Results
Random Forest 0.822 0.823 0.822 0.821
Logistic Regression 0.926 0.926 0.926 0.926
SGDClassifier 0.928 0.930 0.928 0.929
Support Vector Machine 0.939 0.939 0.939 0.939
(a) Stack Overflow
Classifier Accuracy Precision Recall F1 Score
@@
Support Vector Machine 0.999 0.998 0.998 0.998
(b) WWC 2019
@@
Support Vector Machine 0.999 0.999 0.999 0.999
(c) Twitter Trolls
@@
7 Limitations and Scope
While the proposed approach demonstrates strong performance across multiple datasets and tasks, it is important to clarify its scope and limitations. First, the effectiveness of the method depends on the availability and quality of textual properties associated with nodes and relationships. In graphs where textual attributes are sparse, highly noisy, or semantically uninformative, the benefit of text embeddings may be limited. In such cases, structural signals or task-specific features may play a more dominant role.
Second, the approach intentionally does not incorporate explicit structural encoders or message-passing mechanisms. While this design choice enables a lightweight and model-agnostic integration into existing labeled property graph pipelines, it also means that higher-order structural patterns are not directly captured by the embedding process. Tasks that rely heavily on global topology or multi-hop relational reasoning may therefore require complementary structural models or hybrid approaches that combine language-based embeddings with graph-aware encoders.
Finally, the current evaluation focuses on node classification and relation prediction as representative analysis tasks. While these tasks cover common use cases in property graph analytics, further evaluation on additional tasks - such as clustering, anomaly detection, or semantic search - would provide a more comprehensive assessment of the approach’s generality. Despite these limitations, the proposed framework offers a practical and extensible foundation for incorporating textual semantics into labeled property graph analysis without modifying existing graph structures or workflows.
@@ 8 Next Steps
Several directions emerge for extending the proposed framework. A natural next step is the integration of textual embeddings with topology-aware graph encoders. While the current approach deliberately separates semantic representation from graph structure, combining language-model embeddings with graph neural networks or other structural encoders could enable joint reasoning over textual content and multi-hop relational patterns. Such hybrid models may be particularly beneficial for tasks that depend on global graph structure or complex relational dependencies.
Another promising direction concerns task-specific adaptation of the embedding space. Although pretrained text embedding models provide strong general-purpose representations, lightweight fine-tuning or adaptation strategies could further improve performance in domain-specific settings. This includes adapting embeddings to specialized vocabularies, domain terminology, or task objectives commonly encountered in institutional or knowledge-centric graphs.
Scalability and dynamic graph support also represent important avenues for future work. Large labeled property graphs are often subject to frequent updates, including the addition of new nodes, relationships, and textual properties. Developing efficient mechanisms for incremental embedding updates and embedding lifecycle management would be essential for supporting real-time or near-real-time analytics.
@@ 9 Conclusions
We presented an approach for enhancing labeled property graph analysis using pretrained text embedding models. By encoding textual properties into semantic vectors, we enable more informed classification and relation prediction without modifying the underlying graph structure. This method is model-agnostic, lightweight, and compatible with existing property graph platforms. Our evaluation demonstrates its broad applicability across domains.
The results highlight the potential of treating textual node and edge attributes as semantically rich signals, supporting more expressive and effective analysis workflows within graph-based systems. This framework shows that semantic text embeddings can serve as drop-in, model-independent enhancements for graph analytics-bridging language understanding and structural reasoning in a unified, efficient manner.
{credits}
9.0.1 Acknowledgements
This manuscript acknowledges the use of ChatGPT [17], powered by the GPT-5.2 language model developed by OpenAI, to improve language clarity, refine sentence structure, and enhance overall writing precision.